# A stochastic volatility approximation for a tick-by-tick price model with mean-field interaction

### v2

@@ 3.3 Convergence of the joint price-volatility process
Compensating as above the Poisson Processes we obtain
@@
$M_{N}^{\Pi}=\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\left[\int_{[0,\sqrt{N}t]\times[0,+\infty)}{\bf 1}_{[0,\lambda_{N}^{+}(s-))}(u)\tilde{n}_{i}^{+}(ds,du)-\int_{[0,\sqrt{N}t]\times[0,+\infty)}{\bf 1}_{[0,\lambda_{N}^{-}(s-))}(u)\tilde{n}_{i}^{-}(ds,du)\right]$
@@
where $M_{N}^{Z}(t)=\frac{1-\gamma}{2}M_{N}^{\Pi}(t)$. Expanding $f$ as above we obtain
@@
where $Q_{N}(t)$ comes from the cubic part of the expansion, and $Q_{N}(t\wedge\tau_{h}^{N})$ converges uniformly to zero as $N\rightarrow+\infty$. We can now use (55) to eliminate the divergent term from (54).
@@
We are now ready to use Theorem A.1 for the two dimensional process $(\Pi_{N},Y_{N})$. The convergence of $Y_{N}$ has been proved in Proposition 3.5. The initial condition for the limit process $\pi$ follows from (56) and Lemma 3.4, which guarantees that the term $\frac{2f^{\prime}(0)}{\alpha-(1-\gamma)f^{\prime}(0)}Z_{N}(t)$ is negligible for $t\in(0,T]$. Note that it is not so for $t=0$ as, by (22), $Z_{N}(0)$ does not go to zero unless $a^{+}=a^{-}$; for this reason the convergence we prove is in $(0,T]$.
@@
which implies (9).
@@
Step 2: definition of the volatility process and collapsing of $Z_{N}$ (Lemmas 3.2, 3.3 and 3.4)
@@
$\lambda_{N}^{+}(t)=f\left(\frac{1}{\sqrt{N}}\frac{1+\beta_{i}\gamma_{i}}{1+\overline{\beta\gamma}}Y_{N}(t/\sqrt{N})+\frac{2}{\sqrt{N}}\frac{\overline{\beta\gamma}-\beta_{i}\gamma_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}Z_{N}(t/\sqrt{N})\right)$
@@
$\lambda_{N}^{-}(t)=f\left(\frac{1}{\sqrt{N}}\frac{1+\beta_{i}\gamma_{i}}{1+\overline{\beta\gamma}}Y_{N}(t/\sqrt{N})-\frac{2}{\sqrt{N}}\frac{\gamma_{i}\overline{\beta\gamma}-\overline{\gamma}-\beta_{i}\gamma_{i}\overline{\gamma}+\gamma_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}Z_{N}(t/\sqrt{N})\right)$
@@
$\begin{split}\tilde{Z}_{N}^{4}(t)&=E_{N}^{4}+4F_{N}\int_{0}^{t}\tilde{Z}^{2}_{N}(s)ds-4\alpha\sqrt{N}\int_{0}^{t}\tilde{Z}_{N}^{4}(s)ds\\ &+\sqrt{N}\int_{0}^{t}\left[\frac{(1-\overline{\gamma})^{4}}{16N}+\frac{(1-\overline{\gamma})^{3}}{2N^{\frac{3}{2}}}\tilde{Z}_{N}\left(s\right)+3\frac{(1-\overline{\gamma})^{2}}{2N}\tilde{Z}^{2}_{N}\left(s\right)+2\frac{(1-\overline{\gamma})}{\sqrt{N}}\tilde{Z}^{3}_{N}\left(s\right)\right]\\ &\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \sum_{i=1}^{N}f\left(\frac{1}{\sqrt{N}}\frac{1+\beta_{i}\gamma_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)+\frac{2}{\sqrt{N}}\frac{\overline{\beta\gamma}-\beta_{i}\gamma_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)ds\\ &+\sqrt{N}\int_{0}^{t}\left[\frac{(1-\overline{\gamma})^{4}}{16N}-\frac{(1-\overline{\gamma})^{3}}{2N^{\frac{3}{2}}}\tilde{Z}_{N}\left(s\right)+3\frac{(1-\overline{\gamma})^{2}}{2N}\tilde{Z}^{2}_{N}\left(s\right)-2\frac{(1-\overline{\gamma})}{\sqrt{N}}\tilde{Z}^{3}_{N}\left(s\right)\right]\\ &\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \sum_{i=1}^{N}f\left(\frac{1}{\sqrt{N}}\frac{1+\beta_{i}\gamma_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)-\frac{2}{\sqrt{N}}\frac{\gamma_{i}\overline{\beta\gamma}-\overline{\gamma}-\beta_{i}\gamma_{i}\overline{\gamma}+\gamma_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)ds\\ &+M_{N}^{\tilde{Z}^{4}}(t).\end{split}$ (60)
By using the fact that $f$ is Lipschitz and concave, as in (29) and (30), we have
$\sum_{i=1}^{N}\left[f\left(\frac{1}{\sqrt{N}}\frac{1+\beta_{i}\gamma_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)+\frac{2}{\sqrt{N}}\frac{\overline{\beta\gamma}-\beta_{i}\gamma_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)\right.\\ +\left.f\left(\frac{1}{\sqrt{N}}\frac{1+\beta_{i}\gamma_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)-\frac{2}{\sqrt{N}}\frac{\gamma_{i}\overline{\beta\gamma}-\overline{\gamma}-\beta_{i}\gamma_{i}\overline{\gamma}+\gamma_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)\right]\\ \leq\mbox{constant}\cdot\sqrt{N}(\tilde{Y}_{N}(s)+|\tilde{Z}_{N}(s)|)$
@@
$\sum_{i=1}^{N}\left|f\left(\frac{1}{\sqrt{N}}\frac{1+\beta_{i}\gamma_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)+\frac{2}{\sqrt{N}}\frac{\overline{\beta\gamma}-\beta_{i}\gamma_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)\right.\\ -\left.f\left(\frac{1}{\sqrt{N}}\frac{1+\beta_{i}\gamma_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)-\frac{2}{\sqrt{N}}\frac{\gamma_{i}\overline{\beta\gamma}-\overline{\gamma}-\beta_{i}\gamma_{i}\overline{\gamma}+\gamma_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)\right|\\ \leq 2f^{\prime}(0)\cdot\sqrt{N}|\tilde{Z}_{N}(s)|,$
where in this second inequality the special way in which the parameters appear matters. Using this, inequality (32) can be derived and Lemma 3.3 follows for this inhomogeneous model. The proof of Lemma 3.4, which uses a semimartingale representation for $\tilde{Z}_{N}^{2}$, follows along the same lines, and we omit the details.
@@
$\begin{split}Y_{N}(t)&=C_{N}+D_{N}t-\alpha\sqrt{N}\int_{0}^{t}Y_{N}(s)ds\\ &\penalty 10000\ \penalty 10000\ \penalty 10000\ +\frac{\overline{\gamma}(1+\overline{\beta\gamma})}{\overline{\gamma}+\overline{\beta\gamma}}\sum_{i=1}^{N}\int_{0}^{t}f\left(\frac{1}{\sqrt{N}}\frac{1+\beta_{i}\gamma_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)+\frac{2}{\sqrt{N}}\frac{\overline{\beta\gamma}-\beta_{i}\gamma_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)ds\\ &\penalty 10000\ \penalty 10000\ \penalty 10000\ +\frac{\overline{\beta\gamma}(1+\overline{\beta\gamma})}{\overline{\gamma}+\overline{\beta\gamma}}\sum_{i=1}^{N}\int_{0}^{t}f\left(\frac{1}{\sqrt{N}}\frac{1+\beta_{i}\gamma_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)-\frac{2}{\sqrt{N}}\frac{\gamma_{i}\overline{\beta\gamma}-\overline{\gamma}-\beta_{i}\gamma_{i}\overline{\gamma}+\gamma_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)ds\\ &\penalty 10000\ \penalty 10000\ \penalty 10000\ +M_{N}^{Y}(t),\end{split}$ (61)
@@
$\begin{split}\left(\frac{\overline{\gamma}(1+\overline{\beta\gamma})}{\overline{\gamma}+\overline{\beta\gamma}}\right)^{2}\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\int_{0}^{t}f\left(\frac{1}{\sqrt{N}}\frac{1+\beta_{i}\gamma_{i}}{1+\overline{\beta\gamma}}Y_{N}(s)+\frac{2}{\sqrt{N}}\frac{\overline{\beta\gamma}-\beta_{i}\gamma_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}Z_{N}(s)\right)ds\\ +\left(\frac{\overline{\beta\gamma}(1+\overline{\beta\gamma})}{\overline{\gamma}+\overline{\beta\gamma}}\right)^{2}\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\int_{0}^{t}f\left(\frac{1}{\sqrt{N}}\frac{1+\beta_{i}\gamma_{i}}{1+\overline{\beta\gamma}}Y_{N}(s)-\frac{2}{\sqrt{N}}\frac{\gamma_{i}\overline{\beta\gamma}-\overline{\gamma}-\beta_{i}\gamma_{i}\overline{\gamma}+\gamma_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}Z_{N}(s)\right)ds.\end{split}$
@@
$\begin{split}\Pi_{N}(t)&=\sum_{i=1}^{N}\left[\int_{0}^{t}f\left(\frac{1}{\sqrt{N}}\frac{1+\beta_{i}\gamma_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)+\frac{2}{\sqrt{N}}\frac{\overline{\beta\gamma}-\beta_{i}\gamma_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)ds\right.\\ &-\left.\int_{0}^{t}f\left(\frac{1}{\sqrt{N}}\frac{1+\beta_{i}\gamma_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)-\frac{2}{\sqrt{N}}\frac{\gamma_{i}\overline{\beta\gamma}-\overline{\gamma}-\beta_{i}\gamma_{i}\overline{\gamma}+\gamma_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)ds\right]\\ &+M_{N}^{\Pi}(t)\end{split}$ (62)
@@
$\begin{split}\sum_{i=1}^{N}&\left[\int_{0}^{t}f\left(\frac{1}{\sqrt{N}}\frac{1+\beta_{i}\gamma_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)+\frac{2}{\sqrt{N}}\frac{\overline{\beta\gamma}-\beta_{i}\gamma_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)ds\right.\\ &+\left.\int_{0}^{t}f\left(\frac{1}{\sqrt{N}}\frac{1+\beta_{i}\gamma_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)-\frac{2}{\sqrt{N}}\frac{\gamma_{i}\overline{\beta\gamma}-\overline{\gamma}-\beta_{i}\gamma_{i}\overline{\gamma}+\gamma_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)ds\right]\end{split}$ (63)
@@
As before, the criticality condition only depends on dominant terms in the first order expansion of $f$, to which the new extra terms do not contribute. However, the term of the first order expansion $\frac{\kappa}{N}f^{\prime}(0)\sum_{i=1}^{N}Y^{i}_{N}(s)=\kappa f^{\prime}(0)Y_{N}(s)$ has order one, the same as the quadratic term containing $Y^{2}_{N}(s)$. This explains the extra linear term in the drift of the volatility in (17). The rest of the proof requires no change as, again, the derivation of the price dynamics only depends on the linear terms in the expansion of $f$.

### v3

@@ 3.3 Convergence of the joint price-volatility process
Compensating the Poisson Processes as in (26) we obtain
@@
$M_{N}^{\Pi}(t)=\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\left[\int_{[0,\sqrt{N}t]\times[0,+\infty)}{\bf 1}_{[0,\lambda_{N}^{+}(s-))}(u)\tilde{n}_{i}^{+}(ds,du)-\int_{[0,\sqrt{N}t]\times[0,+\infty)}{\bf 1}_{[0,\lambda_{N}^{-}(s-))}(u)\tilde{n}_{i}^{-}(ds,du)\right]$
@@
where $M_{N}^{Z}(t)=\frac{1-\gamma}{2}M_{N}^{\Pi}(t)$. Expanding $f$ as in (47) we obtain
@@
where $Q_{N}(t)$ comes from the cubic part of the expansion; by the same argument used in (54) it is shown that and $Q_{N}(t\wedge\tau_{h}^{N})$ converges uniformly to zero as $N\rightarrow+\infty$. We can now use (59) to eliminate the divergent term from (58).
@@
We are now ready to use Theorem A.1 for the two dimensional process $(\Pi_{N},Y_{N})$. The convergence of $Y_{N}$ has been proved in Proposition 3.6. The initial condition for the limit process $\pi$ follows from (60) and Lemma 3.5, which guarantees that the term $\frac{2f^{\prime}(0)}{\alpha-(1-\gamma)f^{\prime}(0)}Z_{N}(t)$ is negligible for $t\in(0,T]$.
@@
which implies (9) and completes the proof.
We remark that this convergence of the price process $\Pi_{N}(t)$ for $t\in(0,T]$ does not extend to $t=0$; indeed the term $\frac{2f^{\prime}(0)}{\alpha-(1-\gamma)f^{\prime}(0)}Z_{N}(t)$ in (60) does not go to zero unless $a^{+}=a^{-}$.
@@
Step 2: definition of the volatility process and collapsing of $Z_{N}$ (Lemmas 3.3, 3.4 and 3.5)
@@
$\lambda_{N}^{+}(t)=f\left(\frac{1}{\sqrt{N}}\frac{1+\beta^{N}_{i}\gamma^{N}_{i}}{1+\overline{\beta\gamma}}Y_{N}(t/\sqrt{N})+\frac{2}{\sqrt{N}}\frac{\overline{\beta\gamma}-\beta^{N}_{i}\gamma^{N}_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}Z_{N}(t/\sqrt{N})\right)$
@@
$\lambda_{N}^{-}(t)=f\left(\frac{1}{\sqrt{N}}\frac{1+\beta^{N}_{i}\gamma^{N}_{i}}{1+\overline{\beta\gamma}}Y_{N}(t/\sqrt{N})+\frac{2}{\sqrt{N}}\frac{\gamma^{N}_{i}\overline{\beta\gamma}-\overline{\gamma}-\beta^{N}_{i}\gamma^{N}_{i}\overline{\gamma}+\gamma^{N}_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}Z_{N}(t/\sqrt{N})\right)$
@@
$\begin{split}\tilde{Z}_{N}^{4}(t)&=E_{N}^{4}+4F_{N}\int_{0}^{t}\tilde{Z}^{2}_{N}(s)ds-4\alpha\sqrt{N}\int_{0}^{t}\tilde{Z}_{N}^{4}(s)ds\\ &+\sqrt{N}\int_{0}^{t}\left[\frac{(1-\overline{\gamma})^{4}}{16N}+\frac{(1-\overline{\gamma})^{3}}{2N^{\frac{3}{2}}}\tilde{Z}_{N}\left(s\right)+3\frac{(1-\overline{\gamma})^{2}}{2N}\tilde{Z}^{2}_{N}\left(s\right)+2\frac{(1-\overline{\gamma})}{\sqrt{N}}\tilde{Z}^{3}_{N}\left(s\right)\right]\\ &\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \sum_{i=1}^{N}f\left(\frac{1}{\sqrt{N}}\frac{1+\beta^{N}_{i}\gamma^{N}_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)+\frac{2}{\sqrt{N}}\frac{\overline{\beta\gamma}-\beta^{N}_{i}\gamma^{N}_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)ds\\ &+\sqrt{N}\int_{0}^{t}\left[\frac{(1-\overline{\gamma})^{4}}{16N}-\frac{(1-\overline{\gamma})^{3}}{2N^{\frac{3}{2}}}\tilde{Z}_{N}\left(s\right)+3\frac{(1-\overline{\gamma})^{2}}{2N}\tilde{Z}^{2}_{N}\left(s\right)-2\frac{(1-\overline{\gamma})}{\sqrt{N}}\tilde{Z}^{3}_{N}\left(s\right)\right]\\ &\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \sum_{i=1}^{N}f\left(\frac{1}{\sqrt{N}}\frac{1+\beta^{N}_{i}\gamma^{N}_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)+\frac{2}{\sqrt{N}}\frac{\gamma^{N}_{i}\overline{\beta\gamma}-\overline{\gamma}-\beta^{N}_{i}\gamma^{N}_{i}\overline{\gamma}+\gamma^{N}_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)ds\\ &+M_{N}^{\tilde{Z}^{4}}(t).\end{split}$ (64)
By using the fact that $f$ is concave and Lipschitz, as in (31) and (32), we have
$\sum_{i=1}^{N}\left[f\left(\frac{1}{\sqrt{N}}\frac{1+\beta^{N}_{i}\gamma^{N}_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)+\frac{2}{\sqrt{N}}\frac{\overline{\beta\gamma}-\beta^{N}_{i}\gamma^{N}_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)\right.\\ +\left.f\left(\frac{1}{\sqrt{N}}\frac{1+\beta^{N}_{i}\gamma^{N}_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)+\frac{2}{\sqrt{N}}\frac{\gamma_{i}\overline{\beta\gamma}-\overline{\gamma}-\beta^{N}_{i}\gamma^{N}_{i}\overline{\gamma}+\gamma^{N}_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)\right]\\ \leq f^{\prime}(0)\sum_{i=1}^{N}\left[\left(\frac{1}{\sqrt{N}}\frac{1+\beta^{N}_{i}\gamma^{N}_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)+\frac{2}{\sqrt{N}}\frac{\overline{\beta\gamma}-\beta^{N}_{i}\gamma^{N}_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)\right.\\ +\left.\left(\frac{1}{\sqrt{N}}\frac{1+\beta^{N}_{i}\gamma^{N}_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)+\frac{2}{\sqrt{N}}\frac{\gamma_{i}\overline{\beta\gamma}-\overline{\gamma}-\beta^{N}_{i}\gamma^{N}_{i}\overline{\gamma}+\gamma^{N}_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)\right]\\ \leq\mbox{constant}\cdot\sqrt{N}(\tilde{Y}_{N}(s)+|\tilde{Z}_{N}(s)|)$
@@
$\sum_{i=1}^{N}\left|f\left(\frac{1}{\sqrt{N}}\frac{1+\beta^{N}_{i}\gamma^{N}_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)+\frac{2}{\sqrt{N}}\frac{\overline{\beta\gamma}-\beta^{N}_{i}\gamma^{N}_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)\right.\\ -\left.f\left(\frac{1}{\sqrt{N}}\frac{1+\beta^{N}_{i}\gamma^{N}_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)+\frac{2}{\sqrt{N}}\frac{\gamma^{N}_{i}\overline{\beta\gamma}-\overline{\gamma}-\beta^{N}_{i}\gamma^{N}_{i}\overline{\gamma}+\gamma^{N}_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)\right|\\ \leq f^{\prime}(0)\sum_{i=1}^{N}\left|\left(\frac{1}{\sqrt{N}}\frac{1+\beta^{N}_{i}\gamma^{N}_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)+\frac{2}{\sqrt{N}}\frac{\overline{\beta\gamma}-\beta^{N}_{i}\gamma^{N}_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)\right.\\ \left.-\left(\frac{1}{\sqrt{N}}\frac{1+\beta^{N}_{i}\gamma^{N}_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)+\frac{2}{\sqrt{N}}\frac{\gamma^{N}_{i}\overline{\beta\gamma}-\overline{\gamma}-\beta^{N}_{i}\gamma^{N}_{i}\overline{\gamma}+\gamma^{N}_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)\right|\\ \leq 2f^{\prime}(0)\cdot\sqrt{N}|\tilde{Z}_{N}(s)|,$
In the first estimate the generic constant in the last term may depend on the values of all parameters; in the second estimate there are several crucial cancellations to obtain the specific upper bound.
Using these estimates, inequality (34) can be derived and Lemma 3.4 follows for this inhomogeneous model. The proof of Lemma 3.5, which uses a semimartingale representation for $\tilde{Z}_{N}^{2}$, follows along the same lines, and we omit the details.
@@
$\begin{split}Y_{N}(t)&=C_{N}+D_{N}t-\alpha\sqrt{N}\int_{0}^{t}Y_{N}(s)ds\\ &\penalty 10000\ \penalty 10000\ \penalty 10000\ +\frac{\overline{\gamma}(1+\overline{\beta\gamma})}{\overline{\gamma}+\overline{\beta\gamma}}\sum_{i=1}^{N}\int_{0}^{t}f\left(\frac{1}{\sqrt{N}}\frac{1+\beta^{N}_{i}\gamma^{N}_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)+\frac{2}{\sqrt{N}}\frac{\overline{\beta\gamma}-\beta^{N}_{i}\gamma^{N}_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)ds\\ &\penalty 10000\ \penalty 10000\ \penalty 10000\ +\frac{\overline{\beta\gamma}(1+\overline{\beta\gamma})}{\overline{\gamma}+\overline{\beta\gamma}}\sum_{i=1}^{N}\int_{0}^{t}f\left(\frac{1}{\sqrt{N}}\frac{1+\beta^{N}_{i}\gamma^{N}_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)-\frac{2}{\sqrt{N}}\frac{\gamma^{N}_{i}\overline{\beta\gamma}-\overline{\gamma}-\beta^{N}_{i}\gamma^{N}_{i}\overline{\gamma}+\gamma^{N}_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)ds\\ &\penalty 10000\ \penalty 10000\ \penalty 10000\ +M_{N}^{Y}(t),\end{split}$ (65)
@@
$\begin{split}\left(\frac{\overline{\gamma}(1+\overline{\beta\gamma})}{\overline{\gamma}+\overline{\beta\gamma}}\right)^{2}\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\int_{0}^{t}f\left(\frac{1}{\sqrt{N}}\frac{1+\beta^{N}_{i}\gamma^{N}_{i}}{1+\overline{\beta\gamma}}Y_{N}(s)+\frac{2}{\sqrt{N}}\frac{\overline{\beta\gamma}-\beta^{N}_{i}\gamma^{N}_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}Z_{N}(s)\right)ds\\ +\left(\frac{\overline{\beta\gamma}(1+\overline{\beta\gamma})}{\overline{\gamma}+\overline{\beta\gamma}}\right)^{2}\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\int_{0}^{t}f\left(\frac{1}{\sqrt{N}}\frac{1+\beta^{N}_{i}\gamma^{N}_{i}}{1+\overline{\beta\gamma}}Y_{N}(s)-\frac{2}{\sqrt{N}}\frac{\gamma^{N}_{i}\overline{\beta\gamma}-\overline{\gamma}-\beta^{N}_{i}\gamma^{N}_{i}\overline{\gamma}+\gamma^{N}_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}Z_{N}(s)\right)ds.\end{split}$
@@
$\begin{split}\Pi_{N}(t)&=\sum_{i=1}^{N}\left[\int_{0}^{t}f\left(\frac{1}{\sqrt{N}}\frac{1+\beta^{N}_{i}\gamma^{N}_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)+\frac{2}{\sqrt{N}}\frac{\overline{\beta\gamma}-\beta^{N}_{i}\gamma^{N}_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)ds\right.\\ &-\left.\int_{0}^{t}f\left(\frac{1}{\sqrt{N}}\frac{1+\beta^{N}_{i}\gamma^{N}_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)-\frac{2}{\sqrt{N}}\frac{\gamma^{N}_{i}\overline{\beta\gamma}-\overline{\gamma}-\beta^{N}_{i}\gamma^{N}_{i}\overline{\gamma}+\gamma^{N}_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)ds\right]\\ &+M_{N}^{\Pi}(t)\end{split}$ (66)
@@
$\begin{split}\sum_{i=1}^{N}&\left[\int_{0}^{t}f\left(\frac{1}{\sqrt{N}}\frac{1+\beta^{N}_{i}\gamma^{N}_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)+\frac{2}{\sqrt{N}}\frac{\overline{\beta\gamma}-\beta^{N}_{i}\gamma^{N}_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)ds\right.\\ &+\left.\int_{0}^{t}f\left(\frac{1}{\sqrt{N}}\frac{1+\beta^{N}_{i}\gamma^{N}_{i}}{1+\overline{\beta\gamma}}\tilde{Y}_{N}(s)-\frac{2}{\sqrt{N}}\frac{\gamma^{N}_{i}\overline{\beta\gamma}-\overline{\gamma}-\beta^{N}_{i}\gamma^{N}_{i}\overline{\gamma}+\gamma^{N}_{i}\overline{\gamma}}{(1-\overline{\gamma})(\overline{\beta\gamma}+\overline{\gamma})}\tilde{Z}_{N}(s)\right)ds\right]\end{split}$ (67)
@@
As before, criticality condition (16) only depends on dominant terms in the first order expansion of $f$, to which the new extra terms do not contribute. However, the term of the first order expansion $\frac{\kappa}{N}f^{\prime}(0)\sum_{i=1}^{N}Y^{i}_{N}(s)=\kappa f^{\prime}(0)Y_{N}(s)$ has order one, the same as the quadratic term containing $Y^{2}_{N}(s)$. This explains the extra linear term in the drift of the volatility in (17). The rest of the proof requires no change as, again, the derivation of the price dynamics only depends on the linear terms in the expansion of $f$.